OpenStudy (anonymous):

s(x) = 100 + 23x m(x) = 90(1.2x) Complete the table of values. x 100 + 23x 90(1.2x) 0 1 2 3 4 5 Use the table to determine at approximately which point the number of cells will be the same for each organism. Graph the system of equations and show the point of intersection. Explain what the points graphed for each line represent. Explain how you can determine the solutio

OpenStudy (anonymous):

I have the answers i just dont understand the bottom questions

OpenStudy (anonymous):

x 100 + 23x 90(1.2x) 0 100 0 1 123 108 2 146 216 3 169 324 4 192 423 5 215 540 those are the answers i got

OpenStudy (anonymous):

1. x values when m(x) = s(x) , assuming that you have the values and answers on the table. 2. Graph it, and identify the points when line m(x) and line s(x) touch each other. 3. The point represent the number of cells presented in each x. 4. You can determine the solution where you find the intersection point, meaning where they touch. That's the answer. My suggestion, graph it first :)

OpenStudy (anonymous):

for 4. you can also determine by setting m(x) = s(x) 100 + 23x = 90 (1.2x) 100 + 23x = 108 x 100 = 85 x x = 20/17 = 1.176

OpenStudy (anonymous):

ok thanks i'l try that

OpenStudy (anonymous):

@JMark how would i graph this what would i put for my y?

OpenStudy (anonymous):

your y s are the values that you found on your table. m(x) and s(x) have separate y values.

OpenStudy (anonymous):

the ones that you calculated to get

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

ok so i found out they cross at around 146 on the y and like 1 on the x

OpenStudy (anonymous):

Explain how you can determine the solution to the equation 100 + 23x = 90(1.2x) using the graph how would i do this or what would i have to do for this

OpenStudy (anonymous):

Yes. Actually 1.176 = x You can determine the solution by examining the graph. s(x) and m(x) intersect at point (1.176, 146). Because it only intersect at one point, we can determine that the graph has one solution and is independent.

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

you're welcomed :)